OpenStudy (anonymous):

Bernie counted the cars and motorcycles that passed his house in 45 minutes. He counted 26 vehicles and 88 tires. How many motorcycles passed his house?

OpenStudy (turingtest):

let x=#cars y=# of motorcycles "He counted 26 vehicles" means x+y=26 "and 88 tires" means 4x+2y=88 so solve the system x + y=26 4x+2y=88 for y and you have your answer